11 11 TEX L^{A}TEX

Process Monitor HOW-TO for Linux

Al Dev (Alavoor Vasudevan) alavoor@yahoo.com v3.0, 3 December 1999

This document describes how to monitor Linux/Unix processes and to re-start them automatically if they die without any manual intervention. This document also has URLs for "Unix Processes" FAQs.

Contents

1 Linux or Unix Processes 1

2 Unix/Linux command - procautostart 1

3 File procautostart.cpp 2

4 File debug.cpp 13

5 File debug.h 14

6 Makefile 14

7 Testing the program - monitor_test 15

8 Other Monitoring Tools 15

8.1 OpenSource Monitoring Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

8.2 Monitoring Tool - "daemontools" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

8.3 Commercial Monitoring Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

1 Linux or Unix Processes

Processes are the "heart" of the Linux/Unix processes. It is very important to monitor the application processes to ensure 100% availability and reliability of the computer system. For example, processes of databases, web-server etc.. need to be up and running 24 hours a day and 365 days a year. Use the tools described in this document to the monitor important application processes.

Also see the following related topics on Linux/Unix processes.

• Unix Programming FAQ - Chapter 1 Unix Processes <http://www.erlenstar.demon.co.uk/unix/

     faq_toc.html>
   • Other FAQs on Unix are at <http://www.erlenstar.demon.co.uk/unix/>

2 Unix/Linux command - procautostart

Use the program procautostart (say "Prok-Auto-Start" or Process AutoStart) to monitor and automatically re-start any Unix/Linux process if they die. The program listing is given in following sections in this document.

procautostart -n < delay_seconds > -c "< command_line >" nohup &

This starts the unix process procautostart and also command_line process. The procautostart process will re-start command_line process if it dies. The -n option is the time delay in seconds before procautostart checks the running process started by command_line. It is advisable to start the procautostart as background process with no-hangup using "nohup &". See ’man nohup’.

The procautostart is written in "C" so that it is very fast and efficient, since the program is called every n seconds. Amount of resources consumed by procautostart is very minute.

For example -

        procautostart -n 12 -c "monitor_test -d $HOME  -a dummy_arg " nohup &

Here procautostart will be checking the process monitor_test every 12 seconds.

The program will output log files in ’mon’ sub-directory which has datetime stamp of when the processes died and re-started. These files gives info on how often the processes are dying.

You can also use micro-seconds option ’-m’ or nano-seconds option ’-o’, edit the source code file procautostart.cpp and uncomment appropriate lines.

3 File procautostart.cpp

// From your browser save this file as text-file named as ’procautostart.cpp’.

//
// Program to monitor the unix processes
// and automatically re-start them if they die
//
#include <stdio.h>
#include <strings.h> // C strings
#include <unistd.h> // for getopt
#include <alloc.h> // for free
#include <errno.h> // for kill() - error numbers command
extern int errno;
#ifdef Linux
#include <asm/errno.h> // for kill() - error numbers command
#endif
#include <sys/types.h> // for kill() command
#include <signal.h> // for kill() command
#include <sys/wait.h> // for wait()
#include <stdlib.h> // for setenv()
#include <time.h> // for strftime()
#include <libgen.h> // for basename()
#include "debug.h"
#define BUFF_HUN        100
#define BUFF_THOU       1024
#define PR_INIT_VAL     -10
#define WAIT_FOR_SYS    5  // wait for process to start up
#define DEF_SL_SECS     6  // default sleep time
#define SAFE_MEM        10  // to avoid any possible memory leaks
#define LOG_NO          false  // do not output to logfile
#define LOG_YES         true  // do output to logfile
#define STD_ERR_NO      false  // do not print to std err
#define STD_ERR_YES     true  // do print to std err
#define DATE_NO         false  // do not print date
#define DATE_YES        true  // do print date
int start_process(char *commandline, char *args[], char **envp, pid_t proc_pid);
int fork2(pid_t parent_pid, unsigned long tsecs);
inline void error_msg(char *mesg_out, char *lg_file, bool pr_lg, bool std_err, bool pr_dt);

//////////////////////////////////////////////
// To test this program use --
// procautostart -n 5 -c 'monitor_test dummy1 -a dummy2 -b dummy3  ' &
//////////////////////////////////////////////
int main(int argc, char **argv, char **envp)
{
unsigned long   sleep_sec, sleep_micro, sleep_nano;
int     ch;
pid_t   proc_pid;
int pr_no = PR_INIT_VAL;
char mon_log[40];
char *pr_name = NULL, *cmd_line = NULL, **cmdargs = NULL;
// you can turn on debug by editing Makefile and put -DDEBUG in gcc
debug_("test debug", "this line");
debug_("argc", argc);
// Use getpid() - man 2 getpid()
proc_pid = getpid(); // get the Process ID of procautostart
debug_("PID proc_pid", (int) proc_pid);
// Create directory to hold log, temp files
system("mkdir mon 1>/dev/null 2>/dev/null");
sleep_sec = DEF_SL_SECS ; // default sleep time
sleep_micro = 0; // default micro-sleep time
sleep_nano = 0; // default nano-sleep time
optarg = cmd_line = NULL;
while ((ch = getopt(argc, argv, "n:m:o:h:c:")) != -1) // needs trailing colon :
{
        switch (ch)
        {
                case 'n':
                        debug_("scanned option n ", optarg);
                        sleep_sec = atoi(optarg);
                        debug_("sleep_sec", sleep_sec);
                        break;
                case 'm':
                        debug_("scanned option m ", optarg);
                        sleep_micro = atoi(optarg);
                        debug_("sleep_micro", sleep_micro);
                        break;
                case 'o':
                        debug_("scanned option o ", optarg);
                        sleep_nano = atoi(optarg);
                        debug_("sleep_nano", sleep_nano);

break;
case 'c':
debug_("scanned option c ", optarg);
cmd_line = strdup(optarg); // does auto-malloc here
debug_("cmd_line", cmd_line);
break;
case 'h':
debug_("scanned option h ", optarg);
fprintf(stderr, "\nUsage : %s -n <sleep> -m <microsecond> -o <nanosecond> -c '<command>'\n", argv[0]);
exit(-1);
break;
default:
debug_("ch", "default");
fprintf(stderr, "\nUsage : %s -n <sleep> -m <microsecond> -o <nanosecond> -c '<command>'\n", argv[0]);
exit(-1);
break;
}
}
if (cmd_line == NULL)
{
fprintf(stderr, "\ncmd_line is NULL");
fprintf(stderr, "\nUsage : %s -n <sleep> -m <microsecond> -o <nanosecond> -c '<command>'\n", argv[0]);
exit(-1);
}
else
{
// trim the trailing blanks -- otherwise problem in grep command
int tmpii = strlen(cmd_line);
for (int tmpjj = tmpii; tmpjj > -1; tmpjj--)
{
if (cmd_line[tmpjj] == ' ')
cmd_line[tmpjj] = '\0';
else
if (cmd_line[tmpjj] == '&')  // discards amp-and .. we will be appending later
cmd_line[tmpjj] = '\0';
else
if (cmd_line[tmpjj] == '\0')
continue;
else
{
if (cmd_line[tmpjj] == '&') // Discard trailing & in command line
        cmd_line[tmpjj] = '\0';
break;
}
}
debug_("cmd_line", cmd_line);

}
//argv0 = (char *) strdup(argv[0]);
//debug_("argv0", argv0);
// Start the process
{
// Find the command line args
char *aa = strdup(cmd_line), *bb = NULL;
cmdargs = (char **) (malloc(sizeof(char **) + SAFE_MEM));
for (int tmpii = 0; ; tmpii++)
{
        // Allocate more memory ....
        cmdargs = (char **) realloc(cmdargs, (sizeof(char **) * (tmpii+1) + SAFE_MEM)  );
        if (tmpii == 0)
                bb = strtok(aa, " ");
        else
                bb = strtok(NULL, " "); // subsequent calls must have NULL as first arg
        if (bb == NULL)
        {
                cmdargs[tmpii] = bb;
                break;
        }
        else
        {
                // Must malloc with strdup because aa, bb are
                // local vars in local scope!!
                cmdargs[tmpii] = strdup(bb);
        }
        debug_("tmpii", tmpii);
        debug_("cmdargs[tmpii]", (char *) cmdargs[tmpii]);
}
// In case execve you MUST NOT have trailing ampersand & in the command line!!
//pr_no = start_process(cmd_line, NULL, NULL, proc_pid);  // Using execlp ...
pr_no = start_process(cmdargs[0], & cmdargs[0], envp, proc_pid); // Using execve ....
debug_("The child pid", pr_no);
if (pr_no < 0)
{
        fprintf(stderr, "\nFatal Error: Failed to start the process\n");
        exit(-1);
}
sleep(WAIT_FOR_SYS); // wait for the process to come up

// Get process name - only the first word from cmd_line
pr_name = strdup(basename(cmdargs[0])); // process name,  does auto-malloc here
}
// generate log file names
{
char    aa[21];
strncpy(aa, pr_name, 20); aa[20] = '\0';
// Define mon file-names - make it unique with combination of
// process name and process id
sprintf(mon_log, "mon/%s%d.log", aa, (int) proc_pid);
}
// Print out pid to log file
if (pr_no > 0)
{
char aa[200];
sprintf(aa, "Process ID of %s is %d", pr_name, pr_no);
error_msg(aa, mon_log, LOG_YES, STD_ERR_NO, DATE_YES);
}
// monitors the process - restarts if process dies...
bool process_died = false;
char print_log[200];
while (1)  // infinite loop - monitor every 6 seconds
{
//debug_("Monitoring the process now...", ".");
if (kill(pr_no, 0))  // if (kill(pr_no,0) != 0)
{
        debug_("errno from kill() function", errno);
        if (errno == EINVAL)
        {
                process_died = false; // unable to execute kill() - wrong input
                strcpy(print_log, "Error EINVAL: Invalid signal was specified");
                error_msg(print_log, mon_log, LOG_YES, STD_ERR_YES, DATE_YES);
        }
        else
        if (errno == ESRCH )
        {
                // ERSRCH means - No process can be found corresponding to pr_no
                // hence process had died !!
                process_died = true; // No process can be found matching pr_no
                sprintf(print_log,
                "Error ESRCH: No process or process group can be found for %d", pr_no);
                error_msg(print_log, mon_log, LOG_YES, STD_ERR_YES, DATE_YES);

}
else
if (errno == EPERM)
{
process_died = false; // unable to execute kill() - wrong input
strcpy(print_log,
"Error EPERM: The real or saved user ID does not match the real user ID");
error_msg(print_log, mon_log, LOG_YES, STD_ERR_YES, DATE_YES);
}
else
{
process_died = true; // process died!! restart now
debug_("process_die ", "others");
}
if (process_died == true)
{
//
// char respawn[1024];
// strcpy(respawn, cmd_line);
//
// For "C" program use kill(pid_t process, int signal) function.
// #include <signal.h> // See 'man 2 kill'
// Returns 0 on success and -1 with errno set.
//              kill -0 $pid 2>/dev/null || respawn
// To get the exit return status do --
//              kill -0 $pid 2>/dev/null | echo $?
// Return value 0 is success and others mean failure
// Sending 0 does not do anything to target process, but it tests
// whether the process exists. The kill command will set its exit
// status based on this process.
//
// Alternatively, you can use
//              ps -p $pid >/dev/null 2>&1 || respawn
// To get the exit return status do --
//              ps -p $pid >/dev/null 2>&1 | echo $?
// Return value 0 is success and others mean failure
//
// If the process had died, restart and  re-assign the pid to pr_no
// start the process in background ....
// Now re-assign new value of process id to pr_no
if (pr_no > 0 )
        sprintf(print_log, "Fatal Error: Process %s with PID = %d died!!",
                pr_name, pr_no);
else
        sprintf(print_log, "Fatal Error: Process %s is not up!!",

                                pr_name);
                error_msg(print_log, mon_log, LOG_YES, STD_ERR_YES, DATE_YES);
                sprintf(print_log, "Starting process %s", pr_name);
                error_msg(print_log, mon_log, LOG_YES, STD_ERR_NO, DATE_NO);
                //pr_no = start_process(cmd_line, NULL, NULL, proc_pid);  // Using execlp ....
                pr_no = start_process(cmdargs[0], & cmdargs[0], envp, proc_pid); // Using execve ....
                debug_("The child pid", pr_no);
                if (pr_no < 0)
                {
                        sprintf(print_log, "Fatal Error: Failed to start the process");
                        error_msg(print_log, mon_log, LOG_YES, STD_ERR_YES, DATE_YES);
                        exit(-1);
                }
                sleep(WAIT_FOR_SYS); // wait for the process to come up
                sprintf(print_log, "Process ID of %s is %d", pr_name, pr_no);
                error_msg(print_log, mon_log, LOG_YES, STD_ERR_NO, DATE_NO);
        }
}
//debug_("Sleeping now ......", ".");
sleep(sleep_sec);
// Uncomment these to use micro-seconds
// For real-time process control use micro-seconds or nana-seconds sleep functions
// See 'man3 usleep', 'man 2 nanasleep'
// If you do not have usleep() or nanosleep() on your system, use select() or poll()
// specifying no file descriptors to test.
//usleep(sleep_micro);
// To sleep nano-seconds ...  Uncomment these to use nano-seconds
//struct timespec *req = new struct timespec;
//req->tv_sec = 0; // seconds
//req->tv_nsec = sleep_nano; // nanoseconds
//nanosleep( (const struct timespec *)req, NULL);
}
}
inline void error_msg(char *mesg_out, char *lg_file, bool pr_lg, bool std_err, bool pr_dt)
{
if (pr_lg) // (pr_lg == true) output to log file
{
char tmp_msg[BUFF_THOU];
if (pr_dt == true) // print date and message to log file 'lg_file'
{
        sprintf(tmp_msg, "date >> %s; echo '\n%s\n' >> %s\n ",

                lg_file, mesg_out, lg_file);
        system(tmp_msg);
}
else
{
        sprintf(tmp_msg, "echo '\n%s\n' >> %s\n ",
                mesg_out, lg_file);
        system(tmp_msg);
}
}
if (std_err) // (std_err == true) output to standard error
fprintf(stderr, "\n%s\n", mesg_out);
debug_("mesg_out", mesg_out);
}
// start a process and returns PID or -ve value if error
// The main() function has envp arg as in - main(int argc, char *argv[], char **envp)
int start_process(char *commandline, char *args[], char **envp, pid_t parent_pid)
{
int ff;
unsigned long  tsecs;
tsecs = time(NULL); // time in secs since Epoch 1 Jan 1970
debug_("Time tsecs", tsecs);
// Use fork2() instead of fork to avoid zombie child processes
switch (ff = fork2(parent_pid, tsecs))  // fork creates 2 process each executing the following lines
{
case -1:
fprintf(stderr, "\nFatal Error: start_process() - Unable to fork process\n");
_exit(errno);
break;
case 0:  // child process
debug_("\nStarting the start child process\n", " ");
// For child process to ignore the interrupts (i.e. to put
// child process in "background" mode.
// Signals are sent to all processes started from a
// particular terminal. Accordingly, when a program is to be run non-interactively
// (started by &), the shell arranges that the program will ignore interrupts, so
// it won't be stopped by interrupts intended for foreground processes.
// Hence if previous value of signal is not IGN than set it to IGN.
// Note: Signal handlers cannot be set for SIGKILL, SIGSTOP
if (signal(SIGINT, SIG_IGN) == SIG_ERR)
        fprintf(stderr, "\nSignal Error: Not able to set signal to SIGINT\n");

else
if (signal(SIGINT, SIG_IGN) != SIG_IGN) // program already run in background
        signal(SIGINT, SIG_IGN);  // ignore interrupts
if (signal(SIGHUP, SIG_IGN) == SIG_ERR)
        fprintf(stderr, "\nSignal Error: Not able to set signal to SIGHUP\n");
else
if (signal(SIGHUP, SIG_IGN) != SIG_IGN) // program already run in background
        signal(SIGHUP, SIG_IGN);  // ignore hangups
if (signal(SIGQUIT, SIG_IGN) == SIG_ERR)
        fprintf(stderr, "\nSignal Error: Not able to set signal to SIGQUIT\n");
else
if (signal(SIGQUIT, SIG_IGN) != SIG_IGN) // program already run in background
        signal(SIGQUIT, SIG_IGN);  // ignore Quit
if (signal(SIGABRT, SIG_IGN) == SIG_ERR)
        fprintf(stderr, "\nSignal Error: Not able to set signal to SIGABRT\n");
else
if (signal(SIGABRT, SIG_IGN) != SIG_IGN) // program already run in background
        signal(SIGABRT, SIG_IGN);  // ignore ABRT
if (signal(SIGTERM, SIG_IGN) == SIG_ERR)
        fprintf(stderr, "\nSignal Error: Not able to set signal to SIGTERM\n");
else
if (signal(SIGTERM, SIG_IGN) != SIG_IGN) // program already run in background
        signal(SIGTERM, SIG_IGN);  // ignore TERM
// sigtstp - Stop typed at tty. Ignore this so that parent process
// be put in background with CTRL+Z or with SIGSTOP
if (signal(SIGTSTP, SIG_IGN) == SIG_ERR)
        fprintf(stderr, "\nSignal Error: Not able to set signal to SIGTSTP\n");
else
if (signal(SIGTSTP, SIG_IGN) != SIG_IGN) // program already run in background
        signal(SIGTSTP, SIG_IGN);  // ignore TSTP
// You can use debug_generously because they do NOT increase program size!
debug_("before execve commandline", commandline);
debug_("before execve args[0]", args[0]);
debug_("before execve args[1]", args[1]);
debug_("before execve args[2]", args[2]);
debug_("before execve args[3]", args[3]);
debug_("before execve args[4]", args[4]);
debug_("before execve args[5]", args[5]);
debug_("before execve args[6]", args[6]);
debug_("before execve args[7]", args[7]);
execve(commandline, args, envp);

// execlp, execvp does not provide expansion of metacharacters
// like <, >, *, quotes, etc., in argument list. Invoke
// the shell /bin/sh which then does all the work. Construct
// a string 'commandline' that contains the complete command
//execlp("/bin/sh", "sh", "-c", commandline, (char *) 0);  // if success than NEVER returns !!
// If execlp returns than there is some serious error !! And
// executes the following lines below...
fprintf(stderr, "\nFatal Error: Unable to start child process\n");
ff = -2;
exit(127);
break;
default: // parent process
// child pid is ff;
if (ff < 0)
        fprintf(stderr, "\nFatal Error: Problem while starting child process\n");
{
        char    buff[BUFF_HUN];
        FILE    *fp1;
        sprintf(buff, "mon/%d%lu.out", (int) parent_pid, tsecs); // tsecs is unsigned long
        fp1 = fopen(buff, "r");
        if (fp1 != NULL)
        {
                buff[0] = '\0';
                fgets(buff, BUFF_HUN, fp1);
                ff = atoi(buff);
        }
        fclose(fp1);
        debug_("start process(): ff - ", ff);
#ifndef DEBUG
        sprintf(buff, "rm -f mon/%d%lu.out", (int) parent_pid, tsecs);
        system(buff);
#endif // DEBUG
}
// define wait() to put child process in foreground or else put in background
//waitpid(ff, & status, WNOHANG || WUNTRACED);
//waitpid(ff, & status, WUNTRACED);
//wait(& status);
break;
}
return ff;
}

/* fork2() -- like fork, but the new process is immediately orphaned
*            (won't leave a zombie when it exits)
* Returns 1 to the parent, not any meaningful pid.
* The parent cannot wait() for the new process (it's unrelated).
*/
/* This version assumes that you *haven't* caught or ignored SIGCHLD. */
/* If you have, then you should just be using fork() instead anyway.  */
int fork2(pid_t parent_pid, unsigned long tsecs)
{
pid_t mainpid, child_pid = -10;
int status;
    char    buff[BUFF_HUN];
if (!(mainpid = fork()))
{
    switch (child_pid = fork())
    {
      case 0:
                    //child_pid = getpid();
                    //debug_("At case 0 fork2 child_pid : ", child_pid);
                    return 0;
      case -1:
                    _exit(errno);    /* assumes all errnos are <256 */
      default:
                    debug_("fork2 child_pid : ", (int) child_pid);
                    sprintf(buff, "echo %d > mon/%d%lu.out", (int) child_pid, (int) parent_pid, tsecs);
                    system(buff);
                    _exit(0);
    }
}
    //debug_("fork2 pid : ", pid);
if (mainpid < 0 || waitpid(mainpid, & status, 0) < 0)
  return -1;
if (WIFEXITED(status))
  if (WEXITSTATUS(status) == 0)
    return 1;
  else
    errno = WEXITSTATUS(status);
else
  errno = EINTR;  /* well, sort of :-) */
return -1;
}

4 File debug.cpp

// From your browser save this file as text-file named as ’debug.cpp’.

// This file defines the debug_() function which can be used for debugging
// the program. It is similar to "C" assert(). The debug_() is set to void()
// if DEBUG is not defined in Makefile. This way executable size of production
// release is NOT AT ALL effected. Using debug_() very generously has no
// impact on production executable size.
#ifdef DEBUG
#include "debug.h"
// Variable value[] can be char, string, int, unsigned long, float, etc...
void local_dbg(char name[], char value[], char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %s\n", fname, lineno, name, value ); }
void local_dbg(char name[], int value, char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %d\n", fname, lineno, name, value ); }
void local_dbg(char name[], unsigned int value, char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %d\n", fname, lineno, name, value ); }
void local_dbg(char name[], long value, char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %d\n", fname, lineno, name, value ); }
void local_dbg(char name[], unsigned long value, char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %d\n", fname, lineno, name, value ); }
void local_dbg(char name[], short value, char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %d\n", fname, lineno, name, value ); }
void local_dbg(char name[], unsigned short value, char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %d\n", fname, lineno, name, value ); }
void local_dbg(char name[], float value, char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %f\n", fname, lineno, name, value ); }
void local_dbg(char name[], double value, char fname[], int lineno, bool logfile) {
        printf("\nDebug %s Line: %d %s is = %f\n", fname, lineno, name, value ); }
// You add many more here - value can be a class, ENUM, datetime, etc...
#endif // DEBUG

5 File debug.h

// From your browser save this file as text-file named as ’debug.h’.

#ifdef DEBUG
#include <stdio.h>
//#include <strings.h>
//#include <assert.h>  // assert() macro which is also used for debugging
// Debugging code
// Use debug2_to output result to a log file
#define debug_(NM, VL) (void) ( local_dbg(NM, VL, __FILE__, __LINE__) )
#define debug2_(NM, VL, LOG_FILE) (void) ( local_dbg(NM, VL, __FILE__, __LINE__, LOG_FILE) )
void local_dbg(char name[], char value[], char fname[], int lineno, bool logfile= false);
void local_dbg(char name[], int value, char fname[], int lineno, bool logfile= false);
void local_dbg(char name[], unsigned long value, char fname[], int lineno, bool logfile= false);
void local_dbg(char name[], float value, char fname[], int lineno, bool logfile= false);
#else
#define debug_(NM, VL) ((void) 0)
#define debug2_(NM, VL, LOG_FILE) ((void) 0)
#endif // DEBUG

6 Makefile

# From your browser save this file as text-file named as ’Makefile’.

EXE=procautostart
SRCS=procautostart.cpp debug.cpp
OBJS=procautostart.o  debug.o
CXX=gcc
HOSTFLAG=-DLinux
#HOSTFLAG=-DSunOS
# Do not use compiler optimizer -O as this may break the program
# Use debug flag to enable the debug() function. If DEBUG is not
# defined than the function debug() is set to void(), similar
# to assert()
# Use options -Wall (all warning msgs) -O3 (optimization)
#MYCFLAGS=-DDEBUG -g -Wall
MYCFLAGS=-O3 -Wall

all: $(OBJS)
        $(CXX) $(HOSTFLAG) $(MYCFLAGS) $(OBJS) -o $(EXE)
$(OBJS): $(SRCS)
        $(CXX) -c $(HOSTFLAG) $(MYCFLAGS) $(SRCS)
clean:
        rm -f *.o *.log *.log.old *.pid core err a.out afiedt.buf
        rm -f $(EXE)

7 Testing the program - monitor_test

From your browser save this file as text-file named as ’monitor_test’.

Use this program for testing the ’procautostart’ program. For example -

        procautostart -n 12 -c "monitor_test -d $HOME  -a dummy_arg "

Here procautostart will be checking the process monitor_test every 12 seconds.

#!/bin/ksh
# Program to test the procautostart
echo "Started the monitor_test ...."
date > monitor_test.log
while :
do
        date >> monitor_test.log
        sleep 2
done

8 Other Monitoring Tools

8.1 OpenSource Monitoring Tools

On linux systems you can find the following packages. If it is not in the main cdrom than you must check in the contrib cdrom :

• On contrib cdrom daemontools*.rpm

• ’top’ command procps*.rpm

• ’top’ command graphic mode procps-X11*.rpm • ’ktop’ graphic mode ktop*.rpm

• ’gtop’ graphic mode gtop*.rpm

• ’WMMon’ CPU load wmmon*.rpm

• ’wmsysmon’ monitor wmsysmon*.rpm

• ’procmeter’ System activity meter procmeter*.rpm

To use top commands type at unix prompt -

        $ top
        $ ktop
        $ gtop

8.2 Monitoring Tool - "daemontools"

Visit the web site of daemontools at <http://www.pobox.com/~djb/daemontools.html>

To install the daemontools RPM, do -

        # rpm -i /mnt/cdrom/daemontools*.html
        # man supervise

supervise monitors a service. It starts the service and restarts the service if it dies. The companion svc program stops, pauses, or restarts the service on sysadmin request. The svstat program prints a one-line status report. See man page by ’man supervise’

svc - control a supervised service.

svc changes the status of a supervise-monitored service. dir is the same directory used for supervise. You can list several dirs. svc will change the status of each service in turn.

svstat - print the status of a supervised service.

svstat prints the status of a supervise-monitored service. dir is the same directory used for supervise. You can list several dirs. svstat will print the status of each service in turn.

cyclog writes a log to disk. It automatically synchronizes the log every 100KB (by default) to guarantee data integrity after a crash. It automatically rotates the log to keep it below 1MB (by default). If the disk fills up, cyclog pauses and then tries again, without losing any data. See man page by ’man cyclog’

accustamp puts a precise timestamp on each line of input. The timestamp is a numeric TAI timestamp with microsecond precision. The companion tailocal program converts TAI timestamps to local time. See ’man accustamp’

usually watches a log for lines that do not match specified patterns, copying those lines to stderr. The companion errorsto program redirects stderr to a file. See ’man usually’

setuser runs a program under a user’s uid and gid. Unlike su, setuser does not gain privileges; it does not check passwords, and it cannot be run except by root. See ’man setuser’

8.3 Commercial Monitoring Tools

There are commercial monitoring tools available. Check out - • BMC Patrol for Unix/Databases <http://www.bmc.com>

• TIBCO corp’s Hawk for Unix monitoring <http://www.tibco.com>

• LandMark corporation

• Platinum corporation